# Of Mice and Machines: A Comparison of Learning Between Real World Mice and RL Agents

### v1

@@ 8 Discussion
To bring RL agents closer to mice, we introduce two mechanistic innovations: the Post Traumatic Stress Buffer and variance-penalized TD learning. The TISB enhances exploration and strategy diversity, while the variance penalty objective induces information-gathering behaviors reminiscent of biological risk assessment. These mechanisms not only make agent behaviors more mice-like but also advance the interpretability of RL systems by grounding their decision-making processes in biologically plausible behaviors.
@@ Appendix B Experimental Details
All experiments were conducted with an episode length of 300 steps and a total of 100,000 training steps. We use a timestep of 0.25s. This aligns with findings showing that rodents make approximately 2-5 decisions per second during navigation (Resulaj et al., 2009). Through environment setup, we determined that the 0.25s provides an optimal balance between biology realism and learning performance.
@@ B.2 Main Model-Based Method
For VP-TDMPC-2, we modified the buffer and TD-target of the TDMPC-2 agents in their original code. To calculate Q-variance, we first sampled 400 actions uniformly from the action space. We used 5 different ensembled Q-networks, each predicting Q-values for these actions. The variance was then computed across all Q-values produced by these networks.
@@ B.3 Why not SAC and DQN
Notably, with sufficient training, all RL methods ultimately exhibited similar ”reckless” behaviors (lacking waiting periods and showing minimal exploration), confirming that our observations about behavioral gaps are not specific to one algorithm.
@@ Appendix D More Example trajectories
Appendix E Fitting Offline RL On mouse data and motivation to KL divergence
@@ E.1 Behavior Cloning (BC)
Limitations of BC: While BC can effectively mimic certain aspects of the mouse behavior, it struggles with capturing the underlying decision-making strategies, especially in states that are underrepresented or absent in the dataset. This limitation results in poor generalization and inability to solve the task.
@@ E.2 Offline RL Methods
    $\min_{Q}\beta\left(\mathbb{E}_{(s,a)\sim d_{\text{data}}}\left[Q(s,a)\right]-\mathbb{E}_{s\sim d_{\text{data}},a\sim\pi}\left[Q(s,a)\right]\right)+\mathcal{L}_{\text{Bellman}}(Q),$
    where $\beta$ is a regularization parameter, and $\mathcal{L}_{\text{Bellman}}(Q)$ is the Bellman error term.
@@
    Conservative offline model-based policy optimization (COMBO): COMBO (Yu et al., 2021) extends CQL by incorporating a learned model of the environment. It utilizes model-based planning in conjunction with conservative value estimation to improve policy learning. The objective involves optimizing over imaginary rollouts generated by the model while maintaining conservatism:
    $\min_{Q}\beta\left(\mathbb{E}_{(s,a)\sim\hat{d}}\left[Q(s,a)\right]-\mathbb{E}_{s\sim d_{\text{data}},a\sim\pi}\left[Q(s,a)\right]\right)+\mathcal{L}_{\text{Model}}(Q),$
    where $\hat{d}$ represents the distribution induced by the model-generated data, and $\mathcal{L}_{\text{Model}}(Q)$ accounts for model uncertainty.
@@
    Implicit Q-Learning (IQL): IQL (Kostrikov et al., 2021) focuses on recovering optimal actions implicitly by decoupling the value function from policy improvement. Instead of directly minimizing the Bellman error, IQL optimizes:
    $\min_{Q,V}\mathbb{E}_{(s,a)\sim d_{\text{data}}}\left[\left(Q(s,a)-r(s,a)-\gamma V(s^{\prime})\right)^{2}\right],$
    $V(s)=\mathbb{E}_{a\sim\pi}\left[\max\left(Q(s,a)-V(s),0\right)\right].$
    This decoupling allows IQL to better capture the nuanced behaviors present in the dataset.
Figure 17: Offline Training Process for BC, COMBO, CQL, and IQL
Empirical Evaluation: In our predator-prey task, while both CQL and COMBO successfully reproduce mouse trajectories, they fail to emulate the waiting behavior, likely due to an overemphasis on reward optimization leading to more aggressive strategies. In contrast, IQL accurately captures both the waiting actions and the precise trajectory patterns, demonstrating superior adaptability to the intricate behaviors exhibited by mice.
However, the scarcity of mouse trajectories in our dataset prevents us from learning a fully accurate mouse policy. Overcoming this limitation will be a key focus of future work, where we aim to develop more effective behavior‐cloning techniques capable of faithfully reproducing the subtle decision‐making patterns observed in real mice.
To quantitatively assess the discrepancy between the online RL policy and the mouse-derived policy, we employ the Kullback-Leibler (KL) divergence (Hershey & Olsen, 2007) as a metric. For each state $s$ in a representative subset of the state space, the KL divergence between the two policies is defined as:
$KL(\pi_{\text{IQL}}(\cdot|s)\|\pi_{\text{SAC}}(\cdot|s))=\sum_{a}\pi_{\text{IQL}}(a|s)\log\frac{\pi_{\text{IQL}}(a|s)}{\pi_{\text{SAC}}(a|s)}$
Figure 18: KL Divergence between TDMPC-2 and IQL (representing mice policy)
Here, $\pi_{\text{IQL}}(a|s)$ denotes the probability of selecting action $a$ given state $s$ under the IQL policy, and $\pi_{\text{SAC}}(a|s)$ represents the corresponding probability under the SAC policy. By averaging the KL divergence across all sampled states, we obtain an overall measure of the difference between the two policies. Our analysis reveals (Figure 18) that the KL divergence between the online RL policy and the mouse policy is huge, with a value of 6.4386. This significant divergence highlights the limitations of current RL methods in accurately mimicking the complex and nuanced actions exhibited by mice.

### v3

@@ 8 Discussion
These behavioral discrepancies may arise from a fundamental architectural gap. Unlike biological agents that evolve in the context of decisions where the outcome is irreversible (succumbing to a predator, for example), standard RL systems lack an explicit notion of mortality (Ororbia & Friston, 2024). Without permanent failure states, RL agents tend to over-explore and underweight rare but fatal risks, prioritizing immediate goals over the survival-oriented caution observed in biological threat responses. Irreversibility in outcomes has motivated using imagination to inform policy learning in planning algorithms (Racanière et al., 2017), and is a possible basis for the evolution of imagination and planning in the brain (MacIver & Finlay, 2022).
To bring RL agents closer to mice, we introduce two mechanistic innovations: the Trauma-Inspired Safety Buffer and variance-penalized TD learning. The TISB’s dual mechanisms provide key insights: Threat Experience Rebalancing reveals the optimal proportion of negative experiences during early learning, while Threat Experience Amplification enhances wall-following behaviors similar to mice’s defensive strategies. Meanwhile, the variance penalty objective induces information-gathering behaviors reminiscent of biological risk assessment. These mechanisms not only make agent behaviors more mice-like but also advance the interpretability of RL systems by grounding their decision-making processes in biologically plausible behaviors.
@@ Acknowledgments
We gratefully acknowledge the support of the NSF-Simons National Institute for Theory and Mathematics in Biology via grants NSF DMS-2235451 and Simons Foundation MP-TMPS-00005320.
@@ Appendix B Experimental Details
All experiments were conducted with a maximum episode length of 300 steps and a total of 100,000 training steps. We use a timestep of 0.25s. This aligns with findings showing that rodents make approximately 2-5 decisions per second during navigation (Resulaj et al., 2009). Through environment setup, we determined that the 0.25s provides an optimal balance between biology realism and learning performance.
@@ B.2 Main Model-Based Method
For VP-TDMPC-2, we modified the buffer and TD-target of the TDMPC-2 agents in their original code. To calculate Q-variance, we first sampled 400 actions uniformly from the action space. We used an ensemble of 5 different Q-networks, each predicting Q-values for these actions. The variance was then computed across all Q-values produced by these networks.
@@ B.3 Why not SAC and DQN
Notably, with sufficient training, all mentioned RL methods ultimately exhibited similar ”reckless” behaviors (lacking waiting periods and showing minimal exploration), confirming that our observations about behavioral gaps are not specific to one algorithm.
@@ Appendix D More Example trajectories
Appendix E Fitting Offline RL On Mouse Data and Motivation for KL Divergence
To further validate the behavioral differences between mice and RL agents, we employ offline reinforcement learning to directly learn from mouse trajectory data. By training offline RL algorithms on mouse behavioral datasets, we can create ”mouse-like” policies that mimic biological decision-making patterns.
Comparing these learned policies with standard online RL agents provides additional evidence for the fundamental differences in risk assessment and exploration strategies between biological and artificial agents.
@@ E.1 Behavior Cloning (BC)
Limitations of BC: While BC can effectively mimic certain aspects of mouse behavior, it struggles with capturing the underlying decision-making strategies, especially in states that are underrepresented or absent in the dataset. This limitation results in poor generalization and inability to solve the task.
@@ E.2 Offline RL Methods
    $\min_{Q}\;\max_{\mu}\;\alpha\Bigl(\mathbb{E}_{s\sim d_{\text{data}},\,a\sim\pi(a\mid s)}\bigl[Q(s,a)\bigr]\;-\;\mathbb{E}_{(s,a)\sim d_{\text{data}}}\bigl[Q(s,a)\bigr]\Bigr)+\mathcal{L}_{\text{Bellman}}(Q)+\mathcal{R}(\mu),$
    where $\mu$ is an auxiliary action-distribution and $\mathcal{R}(\mu)$ its regularizer.
@@
    Conservative offline model-based policy optimization (COMBO): COMBO (Yu et al., 2021) extends CQL by incorporating a learned model of the environment. It utilizes model-based planning in conjunction with conservative value estimation to improve policy learning:
    $\min_{Q}\alpha\left(\mathbb{E}_{s\sim\hat{d},a\sim\pi(a\mid s)}\left[Q(s,a)\right]-\mathbb{E}_{(s,a)\sim d_{\text{data}}}\left[Q(s,a)\right]\right)+\mathcal{L}_{\text{Bellman}}(Q),$
    where $\hat{d}$ represents the state distribution induced by the learned model.
@@
    Implicit Q-Learning (IQL): IQL (Kostrikov et al., 2021) focuses on recovering optimal actions implicitly by decoupling the value function from policy improvement. It uses expectile regression to learn the value function:
    $\mathcal{L}_{V}(\psi)=\mathbb{E}_{(s,a)\sim d_{\text{data}}}\left[L_{2}^{\tau}\left(Q_{\theta}(s,a)-V_{\psi}(s)\right)\right],$
    $\mathcal{L}_{Q}(\theta)=\mathbb{E}_{(s,a,r,s^{\prime})\sim d_{\text{data}}}\left[\left(Q_{\theta}(s,a)-r-\gamma V_{\psi}(s^{\prime})\right)^{2}\right],$
    where $L_{2}^{\tau}$ is the asymmetric squared loss with expectile $\tau$.
Empirical Evaluation
In our predator-prey task, we evaluate the performance of different offline reinforcement learning methods in reproducing mouse behaviors. While both CQL and COMBO successfully reproduce the basic mouse trajectories, they fail to capture the nuanced waiting behavior observed in real mice. This limitation appears to stem from their overemphasis on reward optimization, which leads to overly aggressive strategies that deviate from the natural behavioral patterns.
In contrast, IQL demonstrates superior performance by accurately capturing both the waiting actions and the precise trajectory patterns exhibited by mice. This success highlights IQL’s enhanced adaptability to complex behavioral dynamics and its ability to balance between reward optimization and behavioral fidelity.
Despite these promising results, the limited availability of mouse trajectory data in our dataset constrains our ability to learn a fully accurate mouse policy. This data scarcity represents a fundamental challenge that must be addressed in future work. We anticipate that developing more sophisticated behavior-cloning techniques will be essential for faithfully reproducing the subtle decision-making patterns observed in real mice.
To provide a quantitative assessment of the behavioral differences, we measure the discrepancy between the online RL policy and the mouse-derived policy using Kullback-Leibler (KL) divergence (Hershey & Olsen, 2007). For each state $s$ in a representative subset of the state space, the KL divergence between the two policies is defined as:
$KL(\pi_{\text{IQL}}(\cdot|s)\|\pi_{\text{RL}}(\cdot|s))=\sum_{a}\pi_{\text{IQL}}(a|s)\log\frac{\pi_{\text{IQL}}(a|s)}{\pi_{\text{RL}}(a|s)}$
where $\pi_{\text{IQL}}(a|s)$ denotes the probability of selecting action $a$ given state $s$ under the IQL policy (representing mouse behavior), and $\pi_{\text{RL}}(a|s)$ represents the corresponding probability under the online RL policy. By averaging the KL divergence across all sampled states, we obtain an overall measure of policy divergence.
Figure 17: KL Divergence between online RL and IQL policies (IQL representing mouse behavior)
Our quantitative analysis reveals a substantial divergence between the online RL policy and the mouse-derived policy, with a KL divergence value of 6.4386 (Figure 17). This significant divergence underscores the considerable gap between current RL approaches and natural mouse behavior, highlighting the complexity of accurately modeling biological decision-making processes. These findings emphasize the need for more sophisticated approaches that can better capture the nuanced behavioral strategies employed by mice in predator-prey scenarios.